### v1

@@ Abstract
The fundamental information-theoretic limits of covert, or low probability of detection/intercept (LPD/LPI), communication have been extensively studied for over a decade, resulting in the square root law (SRL): only $L\sqrt{n}$ covert bits can be reliably transmitted over time-bandwidth product $n$, for constant $L>0$. Transmitting more either results in detection or decoding errors. The SRL imposes significant constraints on hardware realization of mathematically-guaranteed covert communication. Indeed, they preclude using standard link maintenance operations that are taken for granted in non-covert communication. Thus, experimental validation of covert communication is underexplored: to date, only two experimental studies of SRL-based covert communication are available, both focusing on optical channels. Here, we report a demonstration of provably-secure covert radio-frequency (RF) communication using software-defined radios (SDRs). This validates theoretical predictions, opens practical avenues for implementing covert communication systems, and raises further research questions.
@@ I Introduction
Covert, or low probability of detection/intercept (LPD/LPI) communication, allows transmitting messages without alerting an adversary [1, 2, 3, 4]. This mode of information security is highly desirable and contrasts the traditional cryptographic [5] and information-theoretic secrecy [6] methods that prevent adversaries’ access to the transmission’s content, but not its detection. Careful waveform design and spread spectrum techniques are often employed in practice to reduce adversaries’ signal-to-noise ratio (SNR) below the noise floor [7, Pt. 1, Ch. 5]. However, guaranteeing covertness requires intricate mathematical analysis which yields the square root law (SRL): only $B(n)=L\sqrt{n}$ covert bits can be reliably transmitted over $n$ channel uses [1, 2, 3, 4]. The channel-dependent constant $L>0$ is called covert capacity and $n=TW$ is typically the time-bandwidth product for the transmission. Notably, the Shannon capacity [8] of this channel is zero, since $\lim_{n\to\infty}\frac{B(n)}{n}=0$, This is because the adversary’s task in covert communication involves learning just one bit of information (whether transmitter is on or not) versus $\mathcal{O}(n)$ bits of transmitted data in traditional secure communication. Nevertheless, a significant number of such mathematically-guaranteed covert bits can still be transmitted in a finite amount of time.
The discovery of the SRL in [1, 2] resulted in an explosion of research by the communication and information theory communities overviewed in a tutorial [3] and a detailed survey [4]. This included characterization of capacity $L$ for additive white Gaussian noise (AWGN) and discrete memoryless channels (DMCs) [9, 10, 11, 12], covert networks [13, 14, 15, 16, 17, 18], quantum aspects of covert communication [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30], and many other directions. Notwithstanding this progress on fundamental theory, experimental SRL-based covert communication remains underexplored, with only two published works, both focused on optical channels [20, 31]. This paper addresses this gap by providing the first, to our knowledge, experimental validation of the SRL on radio frequency (RF) channels.
@@
We implement and evaluate a covert communication protocol using software-defined radio (SDR). Our implementation uses USRP X310 SDR units deployed on the ORBIT testbed [32, 33], as detailed in Section III-A, enabling controlled and reproducible experiments. We employ quadrature phase-shift keying (QPSK) with a Gaussian pulse shaping filter. This controls temporal and spectral symbol leakage while mitigating timing jitter. We implement an experimental framework based on a synthetic Gaussian noise source within a wired, shielded network that provides environmental control and supports reproducibility.
@@
The strict output power constraint of SRL-based covert communication precludes using standard handshake, carrier synchronization, channel state estimation, and other link maintenance operations that are taken for granted in non-covert communication. This necessitates using precise timing sources, such as atomic clocks. Furthermore, the finite resolution of SDRs’ digital-to-analog and analog-to-digital converters (DACs and ADCs) imposes a lower limit on the amplitude of the transmitted symbol. Thus, to meet the SRL, we must use a “sparse coding” strategy, selecting a random subset of available channel uses that is secretly shared with the receiver in advance. This sparsity, in turn, renders the standard phase synchronization approaches that rely on periodic pilot signal transmission unusable. Therefore, we include a pilot signal with every transmitted symbol, which is done rarely if at all in non-covert communications. This, coupled with precise time synchronization, enables our SDR implementation to transmit a significant number of bits with a covertness guarantee.
The SRL governs, arguably, the worst case scenario. As detailed in Section II, it assumes that the characteristics of the transmitter-adversary channel (noise power variance and transmission loss), the exact time and frequency band of potential transmission, and details of the transceiver system design are known to the adversary. The adversary’s only limitations are: 1) inability to control the random channel noise; and, 2) lack of access to the secret shared between transmitter and receiver prior to the transmission. This is the model implemented in our system. While conservative, it provides a high level of security to unforeseen adversarial technological surprises. Indeed, relaxing these assumptions can result in significant gains in system performance. For example, an adversary’s uncertainty of transmission time/frequency yields a multiplicative improvement to covert capacity $L$ [34, 35, 36], while uncertainty in noise power level can lead to a linear law: $\mathcal{O}(n)$ covert bits reliably transmissible over $n$ channel uses [37, 38]. We defer validation of these to future work. Additionally, our experiment motivates a systematic study of the impact of adversary hardware limitations such as sampling timing jitter and receiver bandwidth constraints. Furthermore, employing more efficient modulation and coding schemes (such as [39]), as well as minimizing the amount of pre-shared secret (see [9]) are avenues for future exploration.
@@
This paper is organized as follows: Section II develops the theoretical foundation for radio-frequency covert communication, beginning with the discrete-time AWGN channel model and deriving a sparse-coded QPSK transmission scheme that satisfies the SRL. Section III details our SDR implementation on the ORBIT testbed, describes the experimental procedure, and presents measurement results that validate the theory. Section IV interprets these findings, highlighting practical design challenges and their implications for real-world systems, and outlines open problems and promising directions for extending this work. Finally, Appendices A-C provide supporting derivations for the covertness criterion, phase-offset estimation, and SNR estimation used in the experiments.
@@ II-A Channel Model
We employ a discrete-time AWGN block channel model in Fig. 1. Alice uses a discrete-time $n_{s}$-length complex-valued pulse shape $\vec{u}(x)\in\mathbb{C}^{n_{s}}$ to modulate data in $x\in\mathcal{X}$, where the modulation constellation $\mathcal{X}$ is defined in Section II-B. This pulse is transmitted over $n_{s}$ channel uses to legitimate receiver Bob and adversary (warden) Willie. Bob receives $h_{b}\vec{u}(x)+\vec{z}^{(b)}$ while Willie observes $h_{w}\vec{u}(x)+\vec{z}^{(w)}$. The respective complex-valued channel gains $h_{b}=a_{b}e^{j\theta_{b}}$ and $h_{w}=a_{w}e^{j\theta_{w}}$ contain attenuation $a_{b},a_{w}>0$ and phase $\theta_{b},\theta_{w}\in[0,2\pi]$ components. We assume channel coherence time longer than the pulse duration, thus, channel gains $h_{b}$ and $h_{w}$ remain constant at least over the $n_{s}$ AWGN channel uses. For each pulse, attenuation is arbitrary, while phase is assumed to be selected uniformly at random from $0,2\pi]$. Each symbol in the transmitted pulse $\vec{u}(x)$ is corrupted independently by AWGN: $\vec{z}^{(b)}$ and $\vec{z}^{(w)}$ are instances of independent and identically distributed (i.i.d.) circularly-symmetric complex Gaussian random vectors $\vec{Z}^{(b)}\sim\mathcal{CN}\left(\vec{0},2\sigma_{b}^{2}\mathbf{I}_{n_{s}}\right)$ and $\vec{Z}^{(w)}\sim\mathcal{CN}\left(\vec{0},2\sigma_{w}^{2}\mathbf{I}_{n_{s}}\right)$, where $\mathbf{I}_{n_{s}}$ is an $n_{s}\times n_{s}$ identity matrix.
The pulse-to-pulse variation in channel gain represents the dynamic operating environment due to the mobility of Alice, Bob, and Willie, as well as the surrounding reflecting objects. Assigning a suitable probability distribution (e.g., Rayleigh, Rice, or Nakagami) to attenuation coefficients $a_{b}$ and $a_{w}$ (along with uniform phase distribution) yields a corresponding block fading model (see [40, Ch. 2]). Here phases $\theta_{b}$ and $\theta_{w}$ are random for each pulse slot, which captures not only environment dynamics but also the drift of independent oscillators. However, we assume for simplicity that attenuation is constant on all channels. This models a scenario where, say, worst-case path-loss figures are employed. We defer the analysis of fully-dynamic channels to future work.
@@ II-B Practical Covert System Design
The square root law (SRL) governs covert communication over the AWGN channel in Fig. 1: one can transmit $B(n)=L\sqrt{n}$ covert bits reliably in $n$ uses of such channel [1, 2], where $L>0$ is the covert capacity (derived in [9, 10]) and $n=TW$ is the transmission time-bandwidth product. These results and much of the follow-up employ a simplified scenario with $n_{s}=1$, $a_{b}=a_{w}=1$, and $\theta_{b}=\theta_{w}=0$. However, practical radio systems require pulse-shaping to efficiently use available bandwidth and to mitigate inter-symbol interference and timing jitter. Furthermore, analysis of the fundamental limits often assumes that Alice can generate inputs at arbitrarily low power. In practice, DACs and ADCs have finite resolution, which results in a lower limit on output signal power. We address this next, and then discuss modulation and pulse shape. Our overall system model is depicted in Fig. 2.
@@ 1 Sparse Coding
Divide $n$ available channel uses into $n_{p}$ pulses, with shape given by the $n_{s}$-length pulse-shape vector $\vec{u}(x)$ discussed next, $x\in\mathcal{X}$. Then, $n=n_{p}n_{s}$, with $n_{p}\in\mathcal{O}(n)$. The minimum output power limit implies the norm $\|\vec{u}(x)\|=c$ for all $x\in\mathcal{X}$. We use sparse coding to ensure covertness: before transmission, Alice and Bob randomly generate and secretly share an $n_{p}$-length sequence $\vec{t}$ of i.i.d. samples from the Bernoulli distribution $p(t_{i})=\{1-\alpha_{n}\text{ if }t_{i}=0;\alpha_{n}\text{ if }t_{i}=1\}$, where the probability of transmitting in a given slot $\alpha_{n}\in\mathcal{O}\left(1/\sqrt{n}\right)$ follows the SRL [1, 2]. The number $n_{t}=\sum_{i=1}^{n_{p}}t_{i}$ of selected pulse slots is a random variable with mean $\alpha_{n}n_{p}$. It is also the length of the transmitted message $\vec{x}_{n_{t}}$, in symbols. Alice is quiet in the slots that are not selected, that is, when $t_{i}=0$, she inputs $\vec{0}$. Denoting by $x=0$ this “innocent” symbol, the transmitted symbol vector $\vec{x}\in\left(\mathcal{X}\cup\{0\}\right)^{n_{p}}$ contains $n_{t}$ symbols from $\mathcal{X}$ and $n_{p}-n_{t}$ innocent symbols. This yields Alice’s sparse-coded covert input $\vec{u}(\vec{x})\in\mathbb{C}^{n}$, containing mostly innocent symbols. We derive $\alpha_{n}$ in Section II-D. We assume that Bob and Alice have a reference clock that allows them to synchronize the start of transmission.
@@ 2 Modulation and Pulse Shape
The channel model in Section II-A includes random pulse-to-pulse phase variation. On-off keying (OOK) modulation is resilient to this, as it uses energy pulses. However, OOK is power-inefficient; matching the per-symbol bit error rate (BER) of binary and quadrature phase-shift keying (BPSK and QPSK) schemes requires transmission of twice the power, which, in turn, doubles Willie’s received SNR and harms covertness. We discuss the relationship between SNR and covert signal construction further in Section II-D. We choose QPSK modulation since it doubles the number of degrees of freedom relative to BPSK, further increasing the achievable covert communication rate.
Phase variations cause errors in phase-keyed systems, which are typically mitigated by periodic pilot signals. However, the SRL does not allow periodic transmissions. Furthermore, the sparse coding scheme results in time between pulses scaling as $\mathcal{O}(\sqrt{n})$ on average, rendering phases uncorrelated for large enough $n$ when channel coherence time is reasonably assumed to be finite. Thus, we include a fixed $n_{s}^{(p)}$-length pilot segment $\vec{c}_{p}$ in every transmitted pulse. For long enough channel coherence time, the pilot segment allows estimation of the phase to correct the $n_{s}^{(q)}$-length QPSK-modulated data segment $\vec{c}_{q}(x)=e^{\frac{j\pi(2x-1)}{4}}\vec{c}_{q}$, $x\in\mathcal{X}$, where $\mathcal{X}=\{1,\ldots,4\}$ and each $x\in\mathcal{X}$ represented by two bits. Further, since the relative phase between pilot and QPSK pulse is what needs to be estimated, Alice randomly selects $\theta$ from a uniform distribution over $0,2\pi)$ and applies to both pilot and QPSK pulse in each transmitted pulse. This improves covertness by ensuring that the relative phase between transmissions is randomized without any additional cost to decoding reliability, since Bob needs to estimate the phase anyway. Hence, the transmitted pulse is given by $\vec{u}(x)=\left(e^{j\theta}\vec{c}_{p},e^{j\theta}\vec{c}_{q}(x)\right)$ as depicted in Fig. 3, with length $n_{s}=n_{s}^{(p)}+n_{s}^{(q)}$ and norm $\|\vec{u}(x)\|\triangleq c$ for all $x\in\mathcal{X}$. Since the pilot requires increased power output per pulse over standard QPSK, this scheme increases Willie’s received SNR based the chosen magnitude of the pilot. Thus, there is a trade-off between phase estimation and covertness of our signal that we optimize in Section III-C.
@@
Sparse coding has negligible inter-symbol interference. Instead covertness demands energy confinement in time and frequency domains. Thus, we shape both $\vec{c}_{p}$ and $\vec{c}_{q}$ by a discrete-time Gaussian envelope:
@@
where $\sigma_{p}$, $\sigma_{q}$, $c_{p}$, and $c_{q}$ are the corresponding pulse width and magnitude parameters. We note that more complicated pulse shapes with lower bandwidth under the same power and time constraints exist, e.g., discrete prolate spheroidal (Slepian) sequences [41, Sec. 3.1] but we defer using them to future work. In Section III-C, we optimize $n_{s}^{(p)}$, $n_{s}^{(q)}$, $c_{p}$, and $c_{q}$. Thus, we choose the largest $\sigma_{k}$ satisfying

### v2

@@ Abstract
The fundamental information-theoretic limits of covert, or low probability of detection/intercept (LPD/LPI), communication have been extensively studied for over a decade, resulting in the square root law (SRL): only $L\sqrt{n}$ covert bits can be reliably transmitted over time-bandwidth product $n$, for constant $L>0$. Transmitting more either results in detection or decoding errors. The SRL imposes significant constraints on the hardware realization of mathematically guaranteed covert communication. These preclude the use of standard link maintenance operations that are taken for granted in non-covert communication. Thus, experimental validation of covert communication is underexplored: to date, only two experimental studies of SRL-based covert communication are available, both focusing on optical channels. Here, we demonstrate provably secure covert radio-frequency (RF) communication using software-defined radios (SDRs). We combine system design, theory, and experiments by 1) developing a sparse-signaling pulse shape to enable covert transmission of data in an environment with potential mobility, 2) proving the covertness of the resulting system, and 3) validating the theoretical predictions by implementing it on SDRs. We uncover and address unique challenges specific to using SDR architecture for covert communications. This opens practical avenues for implementing covert communication systems and raises further research questions.
@@ I Introduction
Covert, or low probability of detection/intercept (LPD/LPI) communication, allows transmitting messages without alerting an adversary [9, 10, 8, 17]. This mode of information security is highly desirable and contrasts with the traditional cryptographic [32] and information-theoretic secrecy [13] methods that prevent adversaries’ access to the transmission’s content, but not its detection. Careful waveform design and spread spectrum techniques are often employed in practice to reduce adversaries’ signal-to-noise ratio (SNR) below the noise floor [37, Pt. 1, Ch. 5]. However, mathematically guaranteeing covertness requires following the square root law (SRL): only $B(n)=L\sqrt{n}$ covert bits can be reliably transmitted over $n$ channel uses [9, 10, 8, 17]. The channel-dependent constant $L>0$ is called covert capacity and $n=TW$ is typically the time-bandwidth product for the transmission. Notably, the Shannon capacity [19] of this channel is zero, since $\lim_{n\to\infty}\frac{B(n)}{n}=0$. This is because the adversary’s task in covert communication involves learning just one bit of information (whether the transmitter is on or not) versus $\mathcal{O}(n)$ bits of transmitted data in traditional secure communication. Nevertheless, a significant number of such mathematically guaranteed covert bits can still be transmitted within a finite time.
The discovery of the SRL in [9, 10] resulted in an explosion of research by the communication and information theory communities, as overviewed in a tutorial [8] and a detailed survey [17]. This included characterization of capacity $L$ for additive white Gaussian noise (AWGN) and discrete memoryless channels (DMCs) [14, 47, 42, 51], covert networks [4, 27, 26, 5, 45, 36, 41, zheng19multiantennacovertcommswirelessnets, liu20iotthz, 18, hieu23ratesplittingcovertmac, kong24covertroutinghetnets, zhao25wardencollusioncovertnets], quantum aspects of covert communication [12, 7, 35, 15, 23, 43, 44, 2, 49, 1, 16, 53], covert communication with unmanned aerial vehicles (UAVs) [jiang21covcommUAVnets, chen21uavjammer, wang23covcommUAV-IRS, tian24uavcovcommasymetricinfo, xu25multiuserUAV, wu25covertUAVsurvey], and many other directions outlined in recent surveys [17, wu25covertUAVsurvey, li25intelligentcovcomm] and references therein.
An adversary lacking knowledge can improve covert communication. For example, an adversary’s uncertainty of transmission time/frequency yields a multiplicative enhancement to covert capacity $L$ [11, 6, 3]. Uncertainty in adversarial channel noise power levels, from, e.g., jammer assistance, can, in theory, lead to a linear law: $\mathcal{O}(n)$ covert bits reliably transmissible over $n$ channel uses [39, 40, 52, chen21uavjammer, zheng21distributedjammingcovcomm]. However, one must approach the assumptions underlying these results with care, e.g., if the underlying channel is continuous-time, transmissions that break the SRL are detectable [24].
On the contrary, the SRL arguably governs the worst case by assuming the adversary knows the transmitter-adversary channel’s characteristics, exact transmission time/frequency, and transceiver design. As detailed in Section II, the adversary’s only limitations are: 1) inability to control the random channel noise; and 2) lack of access to the secret shared between transmitter and receiver prior to the transmission. Therefore, the SRL, while conservative, provides a high level of security against unforeseen adversarial technological surprises. Unfortunately, the strict output power constraint of SRL-based covert communication precludes the use of standard handshakes, carrier synchronization, channel state estimation, and other link-maintenance operations that are taken for granted in non-covert communication. Hence, notwithstanding the progress on fundamental theory, experimental covert communication remains underexplored, with few published works focusing on optical channels [7, 31, Djordjevic25covertSLMatmospheric, Djordjevic25covertSLM].
@@
This paper, to our knowledge, provides the first experimental validation of the SRL on radio frequency (RF) channels. We implement and evaluate a covert communication protocol using software-defined radio (SDR). Section III-A describes our implementation, which uses USRP X310 SDR units on the COSMOS (formerly ORBIT) testbed [33, orbit_grid]. This enables controlled and reproducible experiments. We employ quadrature phase-shift keying (QPSK) with a Gaussian pulse shaping filter. This controls temporal and spectral symbol leakage while mitigating timing jitter. We implement an experimental framework based on a synthetic Gaussian noise source within a wired, shielded network that provides environmental control and supports reproducibility.
@@
Maintaining covert links necessitates using precise timing sources, such as atomic clocks. Furthermore, the finite resolution of modern digital transceivers’ (such as SDRs) digital-to-analog and analog-to-digital converters (DACs and ADCs) imposes a fundamental lower limit on the amplitude of the transmitted symbol. Thus, to meet the SRL, we must use a “sparse coding” strategy, selecting a random subset of available channel uses that is secretly shared with the receiver in advance. This sparsity, in turn, renders the standard phase synchronization approaches that rely on periodic pilot signal transmission unusable. Therefore, we include a pilot signal with every transmitted symbol, which is done rarely if at all in non-covert communications. This, along with precise time synchronization, enables our SDR implementation to transmit a significant number of bits with a guarantee of covertness.
While we consider the worst-case adversary with complete access to the covert communication system’s clock, our experiment motivates a systematic study of the impact of adversary hardware limitations, such as sampling timing jitter, as well as receiver bandwidth constraints. Furthermore, employing more efficient modulation and coding schemes (such as [48]), as well as minimizing the amount of pre-shared secret (see [14]), are avenues for future exploration. Finally, other intermittent communication systems may benefit from our approaches.
@@
This paper is organized as follows: Section II develops the theoretical foundation for radio-frequency covert communication, beginning with the discrete-time AWGN channel model and deriving a sparse-coded QPSK transmission scheme that satisfies the SRL. Section III details our SDR implementation on the COSMOS testbed and describes the experimental procedure, while Section IV presents measurement results that validate the theory. Section V interprets these findings, highlighting practical design challenges and their implications for real-world systems, and outlines open problems and promising directions for extending this work. Finally, Appendices provide supporting derivations for the covertness criterion, phase-offset estimation, and SNR estimation used in the experiments.
@@ II-A Channel Model
As our SDRs are digital and time-synchronized (see Section III-A for details), the discrete-time AWGN block channel model in Fig. 1 describes their input-output relationship in our physical system. A channel use corresponds to a discrete sample of Alice’s input, corrupted by loss and AWGN, as observed by the legitimate receiver Bob and adversary (warden) Willie. Alice uses a discrete-time $n_{s}$-length complex-valued pulse shape $\vec{u}(x)\in\mathbb{C}^{n_{s}}$ to modulate data in $x\in\mathcal{X}$, where the modulation constellation $\mathcal{X}$ is defined in Section II-B. She transmits this pulse over $n_{s}$ channel uses to Bob and Willie. Bob receives $h_{b}\vec{u}(x)+\vec{z}^{(b)}$ while Willie observes $h_{w}\vec{u}(x)+\vec{z}^{(w)}$. The respective complex-valued channel gains $h_{b}=a_{b}e^{j\theta_{b}}$ and $h_{w}=a_{w}e^{j\theta_{w}}$ contain attenuation $a_{b},a_{w}>0$ and phase $\theta_{b},\theta_{w}\in[0,2\pi]$ components. We assume channel coherence time longer than the pulse duration; thus, the channel gains $h_{b}$ and $h_{w}$ are constant for at least $n_{s}$ AWGN channel uses. For each pulse, attenuation is arbitrary, while phase is assumed uniformly random in $0,2\pi]$. Each symbol in the transmitted pulse $\vec{u}(x)$ is corrupted independently by AWGN: $\vec{z}^{(b)}$ and $\vec{z}^{(w)}$ are instances of independent and identically distributed (i.i.d.) circularly-symmetric complex Gaussian random vectors $\vec{Z}^{(b)}\sim\mathcal{CN}\left(\vec{0},2\sigma_{b}^{2}\mathbf{I}_{n_{s}}\right)$ and $\vec{Z}^{(w)}\sim\mathcal{CN}\left(\vec{0},2\sigma_{w}^{2}\mathbf{I}_{n_{s}}\right)$, where $\mathbf{I}_{n_{s}}$ is an $n_{s}\times n_{s}$ identity matrix.
The pulse-to-pulse variation in channel gain models the dynamic operating environment caused by the mobility of Alice, Bob, and Willie, as well as the surrounding reflecting objects. Assigning a suitable probability distribution (e.g., Rayleigh, Rice, or Nakagami) to attenuation coefficients $a_{b}$ and $a_{w}$ (along with uniform phase distribution) yields a corresponding block fading model (see [46, Ch. 2]). Phases $\theta_{b}$ and $\theta_{w}$ are random for each pulse slot, which captures not only environment dynamics but also the drift of independent oscillators. However, for simplicity, we assume that attenuation is constant across all channels. This models a scenario where, say, worst-case path-loss figures are employed. We defer the analysis of fully-dynamic channels to future work.
@@ II-B Practical Covert System Design
The square root law (SRL) governs covert communication over the AWGN channel in Fig. 1: one can transmit $B(n)=L\sqrt{n}$ covert bits reliably in $n$ uses of such channel [9, 10], where $L>0$ is the covert capacity (derived in [14, 47]) and $n=TW$ is the transmission time-bandwidth product. These results and much of the follow-up employ a simplified scenario with $n_{s}=1$, $a_{b}=a_{w}=1$, and $\theta_{b}=\theta_{w}=0$. However, practical radio systems require pulse shaping to efficiently use available bandwidth and mitigate inter-symbol interference and timing jitter. Furthermore, analysis of the fundamental limits often assumes that Alice can generate inputs at arbitrarily low power. In practice, DACs and ADCs have finite resolution, which imposes a lower limit on the output signal power. We address this next, and then discuss modulation and pulse shape. Our overall system model is depicted in Fig. 2.
@@ 1 Sparse Coding
Divide $n$ available channel uses into $n_{p}$ pulses, with shape given by the $n_{s}$-length pulse-shape vector $\vec{u}(x)$ discussed next, $x\in\mathcal{X}$. Then, $n=n_{p}n_{s}$, with $n_{p}\in\mathcal{O}(n)$. The minimum output power limit implies the norm $\|\vec{u}(x)\|=c$ for all $x\in\mathcal{X}$. We use sparse coding to ensure covertness: before transmission, Alice and Bob randomly generate and secretly share an $n_{p}$-length sequence $\vec{t}$ of i.i.d. samples from the Bernoulli distribution $p(t_{i})=\{1-\alpha_{n}\text{ if }t_{i}=0;\alpha_{n}\text{ if }t_{i}=1\}$, where the probability of transmitting in a given slot $\alpha_{n}\in\mathcal{O}\left(1/\sqrt{n}\right)$ follows the SRL [9, 10]. The number $n_{t}=\sum_{i=1}^{n_{p}}t_{i}$ of selected pulse slots is a random variable with mean $\alpha_{n}n_{p}$. It is also the length of the transmitted message $\vec{x}_{n_{t}}$, in symbols. Alice is quiet in the slots that are not selected, that is, when $t_{i}=0$, she inputs $\vec{0}$. We assume that Bob and Alice are synchronized by a common clock. Denoting by $x=0$ this “innocent” symbol, the transmitted symbol vector $\vec{x}\in\left(\mathcal{X}\cup\{0\}\right)^{n_{p}}$ contains $n_{t}$ symbols from $\mathcal{X}$ and $n_{p}-n_{t}$ innocent symbols. This yields Alice’s sparse-coded covert input $\vec{u}(\vec{x})\in\mathbb{C}^{n}$, containing mostly innocent symbols. We derive $\alpha_{n}$ in Section II-D.
@@ 2 Modulation and Pulse Shape
The channel model in Section II-A includes random pulse-to-pulse phase variation. On-off keying (OOK) modulation is resilient to this as it uses energy pulses. However, OOK is power-inefficient; matching the per-symbol bit-error rate (BER) of binary and quadrature phase-shift keying (BPSK and QPSK) schemes requires transmitting twice the power, which, in turn, doubles Willie’s received SNR and harms covertness. We discuss the relationship between SNR and covert signal construction further in Section II-D. We choose QPSK modulation since it doubles the number of degrees of freedom relative to BPSK, further increasing the achievable covert communication rate.
Phase variations cause errors in phase-keyed systems and are typically mitigated using periodic pilot signals; however, the SRL does not permit periodic transmissions. Furthermore, the sparse coding scheme results in the time between pulses scaling as $\mathcal{O}(\sqrt{n})$ on average, rendering the phases uncorrelated for sufficiently large $n$ when the channel coherence time is finite. Thus, we include a fixed $n_{s}^{(p)}$-length strictly-positive pilot segment $\vec{c}_{p}$ in every transmitted pulse. For sufficiently long channel coherence time, the channel induces approximately the same phase rotation on both the pilot and data segments, preserving their relative phase. The pilot segment therefore allows estimation of the phase to correct the $n_{s}^{(q)}$-length QPSK-modulated data segment $\vec{c}_{q}(x)=e^{\frac{j\pi(2x-1)}{4}}\vec{c}_{q}$, $x\in\mathcal{X}$, where $\mathcal{X}=\{1,\ldots,4\}$ and each $x\in\mathcal{X}$ is represented by two bits. Alice additionally applies a random phase $\theta\sim\mathcal{U}[0,2\pi)$ to both pilot and QPSK pulse in each transmitted pulse. This improves covertness by ensuring that the relative phase between transmissions is randomized without any additional cost to decoding reliability, since Bob needs to estimate the phase anyway. Hence, the transmitted pulse is given by $\vec{u}(x)=\left(e^{j\theta}\vec{c}_{p},e^{j\theta}\vec{c}_{q}(x)\right)$ as depicted in Fig. 3, with length $n_{s}=n_{s}^{(p)}+n_{s}^{(q)}$ and norm $\|\vec{u}(x)\|\triangleq c$ for all $x\in\mathcal{X}$. Since the pilot requires additional power, this scheme increases Willie’s received SNR. Thus, there is a trade-off between phase estimation and the covertness of our signal that we optimize in Section III-C.
@@
Sparse coding has negligible inter-symbol interference. Instead, covertness demands energy confinement in time and frequency domains. Thus, we shape both $\vec{c}_{p}$ and $\vec{c}_{q}$ by a discrete-time Gaussian envelope:
@@
where $\sigma_{p}$, $\sigma_{q}$, $c_{p}$, and $c_{q}$ are the corresponding pulse width and magnitude parameters. We note that more complicated pulse shapes with lower bandwidth under the same power and time constraints exist, e.g., discrete prolate spheroidal (Slepian) sequences [38, Sec. 3.1] but we defer using them to future work. In Section III-C, we optimize $n_{s}^{(p)}$, $n_{s}^{(q)}$, $c_{p}$, and $c_{q}$. Thus, we choose the largest $\sigma_{k}$ satisfying